OpenStudy (anonymous):

Write an equation is slope-intercept form for the line that passes through the given point and is perpendicular to the graph of the equation. (5,2),y=1/2x-3

OpenStudy (anonymous):

If it is perpendicular, the slope will be the opposite reciprocal of the given line. Therefore y=(-2)x + b. Then plug in the points that you know. 2=(-2)5+b 2=-10 + b 12=b

OpenStudy (anonymous):

the slope will be the negative reciprocal of 1/2, the equation of the line will be in the same form y=mx+b but you first utilize y-y1=m(x-x1) formula to put the equation of the line into y=mx+b form using the the point given in the question. 5,2. So it becomes y-2=m(negative reciprocal)(x-5).

OpenStudy (anonymous):

then solve for y

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.